how do you find the slope of A(8,2) and B(4,1)?

Answer by pmesler(52)

To find the slope of a line given two points you use the formula
m = y2-y1/x2-x1.

So, given the points A(8,2) and B(4,1) we will say that 8 = x1 and 2 = y1
4 = x2 and 1 = y2.
Then all you do is plug in the values for the formula and you will find the slope.
m = 1-2/4-8 = -1/-4 = 1/4
Therefore the slope is 1/4.